# Uniform Norm Error Estimates for 2D Turning Point Problem

### v1

@@ Abstract
This work presents error analysis for a finite element method applied to a two-dimensional singularly perturbed convection-diffusion turning point problem. Utilizing a layer-adapted Shishkin mesh, we prove uniform convergence in the maximum norm in the coarse layer and x-layer regions. The analysis, critically based on the properties of a discrete Green’s function, guarantees the method’s robustness and accuracy in capturing sharp solution layers. There is no work for the coerce region for two-dimensional singularly perturbed turning point problems, and also, we are getting linear order of convergence better than Stynes’ article.
@@ 1 INTRODUCTION
$\ -\varepsilon\Delta u+\textbf{b}(x,y).\nabla u+c(x,y)u$ $=f(x,y),\quad\text{in}\quad\Omega=(-1,1)\times(-1,1)$ (1.1a)
@@
A particular instance of this general problem (1.1) arises when it forms the focus of our study. The convection vector $b(x,y)$ is specified such that $b_{1}(x,y)=xa(x,y)$ and $b_{2}(x,y)=0\;\forall\;(x,y)\in\Omega$. This specialization leads to the following simplified problem:
$-\varepsilon\Delta u+xa(x,y)u_{x}+c(x,y)u$ $=f(x,y),\quad\text{in}\quad\Omega=(-1,1)\times(-1,1)$ (1.2a)
$u$ $=0\quad\text{on}\quad\Gamma=\partial\Omega.$ (1.2b)
@@
Here, we make a very crucial assumption about $\alpha$ and $\beta$ is that $\alpha>5/2$ and $2/3<\beta\leq 1$. Therefore, $\alpha$ cannot be arbitrarily small. In (1.2), there’s an interior layer at $x=0$. It might not have layers at the side boundaries ($x=-1\;\text{and}\;x=1$). Instead, it only has a parabolic kind of layer at the top ($y=1$) and bottom (y=1y=-1) boundaries.
@@
This paper focuses on the finite element analysis of a two-dimensional singularly perturbed problem with a turning point along the line $x=0$, i.e., a turning line defined on the domain $\times(-1,1)$. The combination of a turning point in a two- or higher-dimensional setting presents a significant challenge for both theoretical analysis and numerical approximation. To address this, we employ a classical finite element method on a properly constructed Shishkin mesh that is adapted to the layer structures induced by the perturbation parameter and the turning point.
@@
Researchers have been actively involved in the progression of finite difference schemes for singularly perturbed turning point problems (SPTPPs). Their work presents both single and multiple interior turning points; for example, see [3, 6, 21, 20, 12, 18, 17, 15, 13]. Berger et al. [6] and Farrell [21] conducted analyses of SPTPPs involving a single interior turning point. Their work confirmed a priori bounds on the analytic solution and its derivatives, which were crucial for demonstrating the uniform convergence of their respective numerical techniques. Specifically, Berger et al. [6] derived an $\varepsilon$-uniform error estimate for a modified El-Mistikawy-Werle finite difference scheme. In parallel, Farrell [21] achieved uniform convergence estimates for a broad category of upwinding-type schemes. In a unique contribution, Natesan and Ramanujam [15] introduced a composite numerical approach for SPTPPs that exhibit twin boundary layers. Their method combines a classical numerical scheme with an exponential fitted-difference scheme to generate an approximate solution. To implement this, the authors divided the problem domain into four distinct sub-regions and applied an asymptotic expansion approximation to solve the problem within each segment. Kadalbajoo and Gupta [20] contributed by deriving asymptotic bounds for the derivatives of the continuous solution to SPTPPs featuring twin boundary layers. These bounds were subsequently used to establish second-order $\varepsilon$-uniform convergence for a B-spline collocation method applied to a piecewise uniform layer-adapted mesh. Natesan et al. [13] proved almost first-order $\varepsilon$-uniform convergence for their proposed scheme, which was based on the classical upwind finite difference method(FDMs) implemented on a Shishkin mesh. Separately, Geng and Qian [12] introduced a stretching variable method and a reproducing kernel method for addressing SPTPPs whose solutions present twin boundary layers. In the work by Becher and Roos [3], the Richardson extrapolation method was applied to an upwind scheme, utilizing a piecewise layer-adapted mesh. This application successfully refined the order of accuracy from $N^{-2}\ln N$ to $N^{-2}(\ln N)^{2}$, where $N$ represents the number of intervals in the spatial direction.
@@
Over the past decades, researchers have increasingly directed their efforts toward developing FEMs for SPTPPs. A key milestone was achieved in 1994 when Sun and Stynes [19] presented and rigorously analyzed piecewise linear Galerkin FEMs. This approximation, tested on diverse piecewise equidistant meshes, was designed for SPPs with interior turning points and solutions featuring interior layers. They successfully established parameter-uniform convergence for their method in the standard $L^{2}$ norm and a weighted energy norm. Building upon this work, Becher [23] in 2016 introduced a higher-order finite element approximation for SPPs with interior turning points. Employing Liseikin’s [25] piecewise uniform meshes, Becher’s contribution was significant in obtaining error bounds that were independent of the perturbation parameter in the energy norm. Research by Becher [4] focused on a one-dimensional SPTPPs that exhibited interior layer characteristics in its solution. The problem was discretized via a standard Galerkin FEMs, then stabilized with SDFEM on piecewise equidistant meshes, yielding parameter-independent error bounds in both the energy and SD-norms. Subsequently, in [2, 14, 9], singularly perturbed boundary turning point problems, both with and without delay, were analyzed. The application of SDFEM in these studies achieved parameter-independent convergence, approaching second order, in the maximum norm. Aasna and Rai [14] further developed the SDFEM framework for time-dependent SPPs with boundary turning points. Their strategy involved employing an $\theta$-scheme on an equidistant time mesh and SDFEM on a spatial Shishkin mesh, successfully establishing parameter-uniform stability and convergence estimates in the maximum norm for their method.
@@
To the best of our knowledge, the pointwise error estimate of the finite element method for two-or higher-dimensional SPTPPs has received considerably less attention in the existing literature. In [7] (1.2) is considered, and a uniform energy norm estimate of order $N^{-1}\log^{2}N$ is derived for the SIPG method on a layer-adapted Shishkin mesh. In [10], Stynes used a Galerkin finite element method on a piecewise equidistant mesh to solve singular Perturbation boundary value problems of convection diffusion type in two dimensions without a turning point. That analysis indicated, uniformly convergent in the perturbation parameter of order $N^{-1}\log N$ in global energy norm and of order $N^{-1/2}\log^{3/2}N$ pointwise near the outflow boundary but not in the coarse region. Key strength is obtained by employing a discrete Green’s function, as it allows us to derive the pointwise error estimates. This observation strongly motivates our present work.
@@
Our research seeks to investigate the pointwise error estimates of the classical finite element method on a Shishkin mesh for the problem of type (1.2). In our article, we implement a classical finite element method on a piecewise uniform mesh(known as Shishkin mesh), but we adopt some specific transition parameter along with xaxisx-axis and along with yaxisy-axis as $\lambda_{x}=min\left(\frac{2\varepsilon}{\alpha}\log\left(\frac{1}{\varepsilon}\right),\frac{1}{2}\right),\;\lambda_{y}=min\left(2\sqrt{\frac{\varepsilon}{\beta}}\log\left(\frac{1}{\varepsilon^{3/2}}\right),\frac{1}{4}\right)$ respectively to solve the convection diffusion boundary value problems with homogeneous Dirichlet boundary conditions with a turning point in the domain $(-1,1)\times(-1,1)$. We establish uniform convergence in the perturbation parameter of order $N^{-1}\log^{3}(1/\varepsilon^{3/2})$ pointwise in the coarse region and $x$-layer region.
@@
This paper proceeds as follows: In Section 1, we establish bounds for the solution of the continuous problem (1.2). In Section 2, we present the classical finite element method, describing its key components such as the finite element space, the chosen basis functions, and the layer-adapted piecewise uniform mesh(known as the Shishkin mesh) with specific mesh transitions. Due to the symmetry along the $y$-axis only need to analyze the problem in the domain $0,1]\times[-1,1]$ along with we decompose our domain into four parts $\Omega_{c},\Omega_{f,x},\Omega_{f,y}$ and $\Omega_{f,xy}$ represents coarse region, $x$-layer region, $y$-layer region and $xy$-layer region respectively. In Section 3, we begin by recalling the discrete Green’s function, which plays a key role in deriving pointwise error estimates. We then proceed to derive $L^{2}$-norm estimates for the discrete Green’s function. A key contribution of this section is a $\varepsilon$-independent estimate for the discrete Green’s function when the source point lies in either the coarse region or the $x$-layer region. This estimate serves as a critical component in the proof of our main theorem. In Section 4, we establish $L^{\infty}$(pointwise) error estimates for the diffusion term. To achieve this, we decompose the solution into four parts: R, $F_{1}$, $F_{2}$, and $F_{12}$ corresponding to no-layer(smooth) region, $x$-layer, $y$-layer, and $xy$-layer regions, respectively. Then, we recall an interpolation theorem from [7], followed by the assumption that source points lie in coarse and $x$-layer regions, and we derived pointwise error estimates for all layer regions. In section 5, following the interpolation framework by Zarin and H.Ross, we develop $L^{\infty}$(pointwise) error estimates for the convection and reaction term. The analysis yields $\varepsilon$-independent error bounds provided the source points reside within the coarse and $x$-layer regions. (pointwise) error estimates for the convection and diffusion terms, following the interpolation framework introduced by Zarin and H. Ross. The analysis yields $\varepsilon$-independent error bounds provided the source point resides within the coarse and x-layer regions. Combining these with estimates for convection, diffusion, and reaction components, we establish the main theorem, which confirms uniform convergence in the perturbation parameter $\varepsilon$ specifically within these regions. Based on the estimates for the convection, diffusion, and reaction terms, we derive the main theorem of this work, which demonstrates uniform convergence in the perturbation parameter $\varepsilon$ of order $N^{-1}\log^{3}(1/\varepsilon^{3/2})$ pointwise specifically with these regions. In section 6, based on the theoretical, we present an example of the convection diffusion turning point problem. The double mesh principle is employed to compute the error estimates and determine the rate of convergence for various values of the perturbation parameter $\varepsilon$ and the mesh discretization parameter N. The pointwise error estimates are presented in Tables 1 and 2 when source points are within smooth and $x$-layer regions, respectively. Table 3 provides the convergence behavior in both subdomains, which is observed to be nearly first order. The numerical solution for Example 1 is depicted in the accompanying figure, thereby demonstrating the effectiveness of the proposed approach. To the best of our knowledge, there has not been any work on the pointwise error estimate for the two or higher-dimensional turning point problems.
@@ 1.1 The continuous problem
The weak formulation of the problem (1.1) is to find an unique $u\in H_{0}^{1}(\Omega)=\{u\in H^{1}(\Omega);u|_{\partial\Omega}=0\}$ such that
@@
Theorem 1
If we assume $c(x,y)-\frac{1}{2}\frac{\partial b_{1}(x,y)}{\partial x}\geq\gamma_{1}>0\;\forall\;(x,y)\in\bar{\Omega}$ then for B(.,.) given in (1.4) the following holds
@@
Singularly perturbed differential equations are typically characterized by a small parameter $\varepsilon<<1$ multiplying some or all of the highest order terms in the differential equation. The solution of such an equation possesses interior or boundary layers or both. These are very thin regions where the derivative of the solution changes very rapidly. Outside these layers, the solution smoothly approximates that of the lower-order reduced problem (obtained by setting $\varepsilon=0$ in (1.2)). In some situations, such as nonlinearities or discontinuous source terms or a sign change of the convection coefficient, interior layers can also be present within the domain. As $\varepsilon\to 0$, the solution’s derivatives within these layers often scale inversely with powers of $\varepsilon$, indicating a significant loss of uniform regularity across the domain. For the problem (1.1), turning points in SPPs can be identified by a sign change in the convective coefficient within the domain. At a turning point, the convective term locally vanishes or becomes very small. This means that diffusion, even if small (due to $\varepsilon$), can locally become relatively more dominant, influencing the solution’s behavior. Both from a theoretical and computational perspective, turning point SPPs pose a greater challenge than standard SPPs. Standard numerical methods designed for convection-dominated problems may struggle to accurately resolve both the boundary and interior layers, often leading to spurious oscillations or significant errors if not specifically adapted (e.g., using fitted meshes).
@@
Theorem 2 [1] Let $f\in C^{3,\beta}(\bar{\Omega})$ where $\beta\in(2/3,1]$ and $f$ satisfy the compatibility condition f(1,1)=f(1,1)=f(1,1)=f(1,1)=0f(-1,-1)=f(1,1)=f(1,-1)=f(-1,1)=0, then the exact solution of given problem can be broken down as $u=R+F_{1}+F_{2}+F_{12}$, where
@@
$\left|\frac{\partial^{i+j}F_{1}(x,y)}{\partial x^{i}\partial y^{j}}\right|$ $\leq C\varepsilon^{-i}e^{-\frac{\alpha x}{\varepsilon}},$
@@
$\left|\frac{\partial^{i+j}F_{12}(x,y)}{\partial x^{i}\partial y^{j}}\right|$ $\leq C\varepsilon^{-i-j/2}e^{-\frac{\alpha x}{\varepsilon}}\left(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\right).$

### v2

# Uniform norm error estimate for rectangular finite element approximation of a 2D turning point problem

@@ Abstract
This work presents error analysis for a finite element method applied to a two-dimensional singularly perturbed convection-diffusion turning point problem. Utilizing a layer-adapted Shishkin mesh, we prove uniform convergence in the maximum norm in the x-layer regions and $\varepsilon$-independent bounds for the coarse region. The analysis, critically based on the properties of a discrete Green’s function, guarantees the method’s robustness and accuracy in capturing sharp solution layers.
@@ 1 INTRODUCTION
$\ -\varepsilon\Delta u+\textbf{b}(x,y).\nabla u+c(x,y)u$ $=f(x,y),\quad\text{in}\quad\Omega=(-1,1)\times(-1,1),$ (1.1a)
@@
A particular case of the general problem (1.1) is of special interest and constitutes the focus of the present study. In particular, the convection vector $b(x,y)$ is specified such that $b_{1}(x,y)=xa(x,y)$ and $b_{2}(x,y)=0\;\forall\;(x,y)\in\Omega$. This specification leads to the following simplified problem:
$-\varepsilon\Delta u+xa(x,y)u_{x}+c(x,y)u$ $=f(x,y),\quad\text{in}\quad\Omega=(-1,1)\times(-1,1),$ (1.2a)
$u$ $=0,\quad\text{on}\quad\Gamma=\partial\Omega.$ (1.2b)
@@
In the resulting problem (1.2), the convection coefficient vanishes along the line $x=0$; therefore, the problem is referred to as a turning-point problem. Here, we make a very crucial assumption about $\alpha$, namely that $\alpha\geq 5/2$. Consequently, $\alpha$ cannot be chosen arbitrarily small. For (1.2), an interior layer occurs at $x=0$. The solution may not exhibit boundary layers at the lateral boundaries ($x=-1\;\text{and}\;x=1$); instead, it displays parabolic-type layers along the top and bottom boundaries ($y=1$) and (y=1y=-1).
@@
This paper focuses on the finite element analysis of a two-dimensional singularly perturbed problem with turning points along the line $x=0$, i.e., a turning line defined on the domain $\times(-1,1)$. The combination of a turning point in a two or higher-dimensional setting presents a significant challenge for both theoretical analysis and numerical approximation. To address this, we employ a conforming quadrilateral finite element on a properly constructed Shishkin mesh that is adapted to the layer structures induced by the perturbation parameter and the turning point.
@@
Researchers have been actively involved in the progression of finite difference schemes for singularly perturbed turning point problems (SPTPPs). Their work presents both single and multiple interior turning points; for example, see [4, 7, 9, 12, 11, 14, 15, 19, 20]. Berger et al.[7] and Farrell [9] conducted analyses of SPTPPs involving a single interior turning point. Their work confirmed a priori bounds on the analytic solution and its derivatives, which were crucial for demonstrating the uniform convergence of their respective numerical scheme. Specifically, Berger et al.[7] derived an $\varepsilon$-uniform error estimate for a modified El-Mistikawy-Werle finite difference scheme. In parallel, Farrell [9] achieved uniform convergence estimates for a broad category of upwinding-type schemes. In a unique contribution, Natesan and Ramanujam [19] introduced a composite numerical approach for SPTPPs that exhibit twin boundary layers. Their method combines a classical numerical scheme with an exponential fitted-difference scheme to generate an approximate solution. To implement this, the authors divided the problem domain into four distinct sub-regions and applied an asymptotic expansion approximation to solve the problem within each segment. Kadalbajoo and Gupta [12] contributed by deriving asymptotic bounds for the derivatives of the continuous solution to SPTPPs featuring twin boundary layers. These bounds were subsequently used to establish second-order $\varepsilon$-uniform convergence for a B-spline collocation method applied to a piecewise uniform layer-adapted mesh. Natesan et al.[20] proved almost first-order $\varepsilon$-uniform convergence for their proposed scheme, which was based on the classical upwind finite difference method (FDMs) implemented on a Shishkin mesh. Separately, Geng and Qian [11] introduced a stretching variable method and a reproducing kernel method for addressing SPTPPs whose solutions present twin boundary layers. In the work by Becher and Roos [4], the Richardson extrapolation method was applied to an upwind scheme, utilizing a piecewise layer-adapted mesh. This application successfully refined the order of accuracy from $N^{-2}\ln N$ to $N^{-2}(\ln N)^{2}$, where $N$ represents the number of intervals in the spatial direction.
@@
Over the past decades, researchers have increasingly directed their efforts toward developing FEMs for SPTPPs. A key milestone was achieved in 1994 when Sun and Stynes [24] presented and rigorously analyzed piecewise linear Galerkin FEMs. This scheme, tested on diverse piecewise equidistant meshes, was designed for SPPs with interior turning points and solutions featuring interior layers. They successfully established parameter-uniform convergence for their method in the standard $L^{2}$-norm and a weighted energy norm. Building upon this work, Becher [5] in 2016 introduced a higher-order finite element approximation for SPPs with interior turning points. Employing Liseikin’s [17] mesh, Becher’s contribution was significant in obtaining error bounds that were independent of the perturbation parameter in the energy norm. Work by Becher [6] focused on a one-dimensional SPTPPs that exhibited interior layer characteristics in its solution. The problem was discretized via a standard Galerkin FEMs, then stabilized with SDFEM on piecewise equidistant meshes, yielding parameter-independent error bounds in both the energy and SD-norms. Subsequently, in [1, 2, 8] singularly perturbed boundary turning point problems, both with and without delay, were analyzed. The application of SDFEM in these studies achieved parameter-independent convergence, approaching second order, in the maximum norm. Aasna and Rai [1] further developed the SDFEM framework for time-dependent SPPs with boundary turning points. Their strategy involved employing an $\theta$-scheme on an equidistant time mesh and SDFEM on a spatial Shishkin mesh, successfully establishing parameter-uniform stability and convergence estimates in the maximum norm for their method. Ranjan [21] considered (1.2), and a uniform energy norm estimate of order $N^{-1}\log^{2}N$ is derived for the SIPG method on a layer-adapted Shishkin mesh.
@@
To the best of our knowledge, the pointwise error estimate of the finite element method for two-or higher-dimensional SPTPPs has received considerably less attention in the existing literature. In [23], Stynes used a Galerkin finite element method on a piecewise equidistant mesh to solve a singularly perturbed boundary value problem of convection diffusion type in two dimensions without a turning point. That analysis established uniform convergence with respect to the perturbation parameter of order $N^{-1}\log N$ in the global energy norm, and of order $N^{-1/2}\log^{3/2}N$ pointwise near the outflow boundary, but not in the coarse region. A key strength of the analysis lies in the use of a discrete Green’s function, which enables the derivation of pointwise error estimates. This observation strongly motivates the present work. In [27], pointwise error estimates based on discrete Green’s functions for the SDFEM were established by Zhang for the convection diffusion problems containing only boundary layers. Though the analysis was based on a uniform mesh. Pointwise energy norm and $L^{2}$-norm estimates on layer-adapted meshes for purely boundary layer and turning point convection diffusion problems were derived in [16] for SDFEM. There, the barrier function technique is used instead of the discrete Green’s function technique.
In this work, we investigate pointwise error estimates for the classical finite element method applied to problems of type (1.2) on a Shishkin mesh. We employ the standard finite element method on a piecewise uniform (Shishkin) mesh, incorporating specific transition parameters along the $x$- and $y$-directions, given respectively by
$\lambda_{x}=\min\left(\frac{2\varepsilon}{\alpha}\log\left(\frac{1}{\varepsilon}\right),\frac{1}{2}\right),\quad\lambda_{y}=\min\left(2\sqrt{\frac{\varepsilon}{\beta}}\log\left(\frac{1}{\varepsilon^{3/2}}\right),\frac{1}{4}\right).$
@@
This framework is used to solve convection-diffusion boundary value problems with homogeneous Dirichlet boundary conditions and a turning point in the domain $(-1,1)\times(-1,1)$. We establish pointwise uniform convergence with respect to the perturbation parameter, achieving an order of $\frac{(\log(1/\varepsilon))^{1/2}}{N^{1/2}}$ in the $x$-layer region. To the best of our knowledge, this is the first work to provide a pointwise norm error estimate for turning point problems employing the classical finite element method on the layer-adapted meshes.
@@
This paper proceeds as follows: In Section 1, we establish bounds for the solution of the continuous problem (1.2). In Section 2, we present the classical finite element method, describing its key components such as the finite element space, the chosen basis functions, and the layer-adapted piecewise uniform mesh (known as the Shishkin mesh) with specific mesh transitions. Due to the symmetry along the $y$-axis only need to analyze the problem in the domain $0,1]\times[-1,1]$ along with we decompose our domain into four parts $\Omega_{c},\;\Omega_{f,x},\;\Omega_{f,y}$ and $\Omega_{f,xy}$ represents coarse region, $x$-layer region, $y$-layer region and $xy$-layer region respectively. In Section 3, we begin by recalling the discrete Green’s function, which plays a key role in deriving pointwise error estimates. Then, we proceed to derive $L^{2}$-norm estimates for the discrete Green’s function. A key contribution of this section is a $\varepsilon$-independent estimate for the discrete Green’s function when the source point lies in either the coarse region or the $x$-layer region. This estimate serves as a critical component in the proof of our main theorem. In Section 4, we establish $L^{\infty}$-error (pointwise) estimates for the diffusion term. To achieve this, we decompose the solution into four parts: R, $F_{1}$, $F_{2}$, and $F_{12}$ corresponding to no-layer (smooth) region, $x$-layer, $y$-layer, and $xy$-layer regions, respectively. Then, we recall an interpolation theorem from [21], followed by the assumption that source points lie in coarse and $x$-layer regions, and we derived pointwise error estimates for all layer regions. In section 5, following the interpolation framework by Zarin and H. Ross, we develop $L^{\infty}$-norm error (pointwise) estimates for the convection and reaction term. The analysis yields $\varepsilon$-independent error bounds provided the source point resides within the coarse and x-layer regions. In section 6, combining estimates for convection, diffusion, and reaction components, we establish the main theorem, which confirms uniform convergence in the perturbation parameter ($\varepsilon$) specifically within these regions. Based on the estimates for the convection, diffusion, and reaction terms, we derive the main theorem of this work, which demonstrates uniform convergence in the perturbation parameter $\frac{(\log(1/\varepsilon))^{1/2}}{N^{1/2}}$ pointwise specifically $(x_{m},y_{n})\in\Omega_{f,x}$ and $\varepsilon$-independent bounds when $(x_{m},y_{n})\in\Omega_{c}$. In section 7, based on the theoretical, we present an example of the convection diffusion turning point problem. The double mesh principle is employed to compute the error estimates and determine the rate of convergence for various values of the perturbation parameter $\varepsilon$ and the mesh discretization parameter N. The pointwise error estimates are presented in Table 1 and Table 2 when source points are within smooth and $x$-layer regions, respectively. Table 3 provides the convergence behavior in both subdomains. The numerical solution for Example 7.1 is depicted in the accompanying figure, thereby demonstrating the effectiveness of the proposed approach. To the best of our knowledge, there has not been any work on the pointwise error estimate for the two or higher-dimensional turning point problems.
@@ 1.1 The continuous problem
The weak formulation for (1.2) is to find an unique $u\in H_{0}^{1}(\Omega)=\{u\in H^{1}(\Omega);u|_{\partial\Omega}=0\}$ such that
@@
Theorem 1.1
[22] If we assume $c(x,y)-\frac{1}{2}\frac{\partial b_{1}(x,y)}{\partial x}\geq\gamma_{1}>0\;\forall\;(x,y)\in\bar{\Omega}$, then for B(.,.) given in (1.4) the following holds
@@
Singularly perturbed differential equations are typically characterized by a small parameter $\varepsilon<<1$ multiplying some or all of the highest order terms in the differential equation. The solution of such an equation possesses interior or boundary layers or both. These are very thin regions where the solution and its derivative change very rapidly. Outside these layers, the solution smoothly approximates that of the lower-order reduced problem (obtained by setting $\varepsilon=0$ in (1.2)). In some situations, such as nonlinearities or discontinuous source terms or a sign change in the convection coefficient, interior layers can also be present within the domain. As $\varepsilon\to 0$, the solution’s derivative within these layers often scales inversely with powers of $\varepsilon$, indicating a significant loss of uniform regularity across the domain. For the problem (1.2), turning points in SPPs can be identified by a sign change in the convective coefficient within the domain. At a turning point, the convective term locally vanishes or becomes very small. This means that diffusion, even if small (due to $\varepsilon$), can locally become relatively more dominant, influencing the solution behavior. From both theoretical and computational perspectives, turning point SPPs pose a greater challenge than standard SPPs. Standard numerical methods designed for convection-dominated problems may struggle to accurately resolve both the boundary and interior layers, often leading to spurious oscillations or significant errors if not specifically adapted (e.g., using fitted meshes).
@@
Theorem 1.2
[18],[22] Let $f\in C^{3,\beta}(\bar{\Omega})$ where $\beta\in[2/3,1]$ and $f$ satisfy the compatibility condition f(1,1)=f(1,1)=f(1,1)=f(1,1)=0f(-1,-1)=f(1,1)=f(1,-1)=f(-1,1)=0, then the exact solution of given problem can be broken down as $u=R+F_{1}+F_{2}+F_{12}$, where
@@
$\left|\frac{\partial^{i+j}F_{1}(x,y)}{\partial x^{i}\partial y^{j}}\right|$ $\leq C\varepsilon^{-i}e^{-\frac{\alpha|x|}{\varepsilon}},$
@@
$\left|\frac{\partial^{i+j}F_{12}(x,y)}{\partial x^{i}\partial y^{j}}\right|$ $\leq C\varepsilon^{-i-j/2}e^{-\frac{\alpha|x|}{\varepsilon}}\left(e^{-\frac{\beta}{\sqrt{\varepsilon}}(1+y)}+e^{-\frac{\beta}{\sqrt{\varepsilon}}(1-y)}\right).$